OpenStudy (anonymous):

the white blood cell concentration in normal blood is approximately 12,00 cells/mm^3. How many white blood cells does a normal adult with 5L of blood have? Express the answer in scientific notation.

OpenStudy (anonymous):

A mm^3 is NOT a liter ... a dm^3 is a liter. (12,000 cells / mm^3) x (100 mm / dm)^3 = (12,000 x (100)^3 cells / dm^3) = 1.2 x 10^10 cells / L That x 5 L = 6.0 x 10^10 cells

OpenStudy (triciaal):

@imsmartygofygirl66 mm^3 is correct in the question! It is the unit of measure. Yes to do this problem we will need to change between liters and milliliters 1000mm = 1L now try again.